OpenStudy (anonymous):

Use the laws of exponents to simplify. Write the answer with positive exponents. (32k^5m^-10)^1/5 I'm not sure how to solve this at all. I'm so frustrated.

OpenStudy (astrophysics):

\[\huge (32k^5m^{-10})^{1/5}\] this right

OpenStudy (anonymous):

Yesh

OpenStudy (astrophysics):

Here are some useful rules you should try out, then if you still have trouble we can go over them \[\huge (x^m)^n = x^{mn}\] \[\huge x^{-n} = \frac{ 1 }{ x^n }\]

OpenStudy (astrophysics):

Note that x is a variable here, you must still apply the 1/5 to the coefficient of 32

OpenStudy (anonymous):

I'm confused with how to use k and m in those equations

OpenStudy (astrophysics):

Can you should me what you have done

OpenStudy (astrophysics):

Apply the first rule to your expression

OpenStudy (anonymous):

I stopped at the coefficient because I didn't know what variables to use

OpenStudy (astrophysics):

\[\huge (32k^5m^{-10})^{1/5} = 32^{1/5}k^{5 \times 1/5}m^{-10 \times 1/5}\]

OpenStudy (astrophysics):

\[\huge 2k^{5/5}m^{-10/5} = 2km^{-2}\] now use the second rule